OpenStudy (anonymous):

Steve has these score on the exams for this semester: 89, 92, and 87. How much will have to make on the fourth exam to have a final average of an A (a 92% or better)?

OpenStudy (anonymous):

solve for x \[\frac{ 89+92+87+x }{ 400 } = .92\]

OpenStudy (anonymous):

\[\frac{ 89+92+87+x }{ 4 }=92\] is how most would set that up, i think

OpenStudy (anonymous):

^_^

OpenStudy (anonymous):

So its %100?

OpenStudy (anonymous):

not quite.. first we need to isolate the 'x' term, so clear fractions would be the first step, so times both sides by the denomenator